The density of gold is 19.3 g/ cm3. What is this value in kilograms per cubic meter?



1
Expert's answer
2022-02-16T08:34:46-0500
"\rm 19.3\: \frac{g}{ cm^3}=19.3\: \frac{1/1000\: kg}{(1/100\: m)^3}=19300\: kg/m^3"